OpenStudy (anonymous):

Find the quotient. (5ab - 10b2 + 15bc) ÷ 5b

OpenStudy (tennistar):

= (5ab - 10b² + 15bc)/5b = (5b[a - 2b + 3c])/5b = a - 2b + 3c Answer: a - 2b + 3c

OpenStudy (tennistar):

you cancel out the 5bs